# Worse than Random? An Embarrassingly Simple Probing Evaluation of Large Multimodal Models in Medical VQA

### v1

@@ 4.1.1 Probing Evaluation with Adversarial Pairs in VQA-RAD
LLaVA-v1-7B 96.61 25.42 (-71.19)
LLaVA-v1.6-7B 77.11 8.47 (-68.64)
LLaVA-Med 81.35 3.38 (-77.97)
MiniGPT-v2 72.03 46.61 (-25.42)
CheXagent 53.38 21.18 (-32.20)
GPT-4V 40.67 33.89    (-6.78)
Gemini Pro 61.86 44.91 (-16.95)
4.1.2 Probing Evaluation with Adversarial Pairs in ProbMed
Table 4 demonstrates the similar significant impact of adversarial pairs in ProbMed. The accuracy of more capable models is generally less affected by the introduction of challenging adversarial pairs. However, even the most robust models experience a minimum drop of 10.52% in accuracy when tested with ProbMed’s challenging questions, with an average decrease of 44.7% across the tested models, highlighting the critical role of probing evaluation in evaluating Med-VQA performance comprehensively.
Table 4: The impact of adversarial pairs on model accuracy in ProbMed. This table illustrates the performance change of models with the introduction of adversarial pairs. The results indicate a substantial decline in accuracy, underlining the robustness of ProbMed in challenging LMMs. The percentage drop in accuracy is noted in parentheses.
Models
Accuracy (%) without
Adversarial Pairs
@@ 4.1.2 Probing Evaluation with Adversarial Pairs in ProbMed
LLaVA-v1-7B 95.40 16.51 (-78.89)
LLaVA-v1.6-7B 83.79 21.97 (-61.82)
LLaVA-Med 71.91 14.77 (-57.14)
MiniGPT-v2 85.30 22.99 (-62.31)
CheXagent 39.55 22.18 (-17.37)
GPT-4V 66.27 55.75 (-10.52)
Gemini Pro 78.36 53.26 (-25.10)
@@ 4.2 How Reliable Are LMMs in Medical Diagnosis?
After correcting model accuracy by introducing adversarial pairs, we continue to address the second research question. We conducted diagnostic probing ranging from general to specialized diagnostic questions using the ProbMed dataset.
4.2.1 Performance across Diagnostic Questions
@@
Table 5 shows the categorical accuracy of different models aggregated among all image types. While GPT-4V and Gemini Pro outperform other models and excel in general tasks such as recognizing image modality and organs, their low performance in specialized tasks like determining the existence of abnormalities and answering fine-grained questions about condition/finding and position highlights a significant gap in their ability to aid in real-life diagnosis.
@@
On more specialized diagnostic questions, even top-performing models like GPT-4V and Gemini Pro performed close to random guessing. Their accuracy in identifying conditions and positions was alarmingly low, underscoring their limitations in handling fine-grained medical inquiries. CheXagent, trained exclusively on chest X-rays, achieved the highest accuracy in determining abnormalities and conditions. However, its performance in general tasks like identifying image modality and organs was lower, highlighting the need for domain-specific training. LLaVA-Med achieves much higher accuracy compared to both LLaVA-v1-7B and LLaVA-v1.6-7B in identifying conditions/finding and their positions but fell short on more general tasks like identifying image modality and organs as well as determining abnormality.
@@
Among the open-sourced general-purpose models, MiniGPT-v2 performs the best, surpassing domain-specific LLaVA-Med in identifying organs and both LLaVA-Med and CheXagent in determining positions of condition/finding without domain-specific training. A more detailed breakdown of the performance of different models on different image types across each question type is available in Appendix A.
Table 5: Categorical and overall accuracy (%) of different models aggregated among all image types in ProbMed. The best result in each question category is in-bold, and the second best is underlined.
General Question Specialized Question
Modality Organ Abnormality Condition/Finding Position
Random Choice 25.00 25.00 50.00 35.67 36.48
LLaVA-v1-7B 25.28 40.53 50.00 0.34 0.10
LLaVA-v1.6-7B 6.77 80.70 46.18 3.57 1.07
LLaVA-Med 5.49 32.98 38.76 20.39 5.37
MiniGPT-v2 3.25 76.29 50.09 15.23 8.05
CheXagent 37.26 33.96 73.32 28.53 7.24
GPT-4V 92.50 71.71 53.30 35.19 22.32
Gemini Pro 96.47 75.69 62.59 27.92 17.17
4.2.2 Error Analysis in Procedural Diagnosis
For models whose accuracy dropped drastically after introducing adversarial pairs, we observed a consistent accuracy pattern much lower than random guess performance for specialized questions. An error analysis focusing on GPT-4V and Gemini Pro across three specialized question types - Abnormality, Condition/Finding, and Position is further conducted. Each accuracy measurement is conditional on the model successfully answering the preceding diagnostic questions, reflecting a procedural diagnosis approach. This analysis reveals both models’ vulnerabilities to hallucination errors, particularly as they progress through the diagnostic procedure, with Gemini Pro being more prone to accepting false conditions and positions.
As shown in Table 6, for Abnormality questions, conditioned on correctly identifying both image modality and organ, GPT-4V’s errors arise from both incorrect answers and its tendency to reject challenging questions, while Gemini Pro attained a slightly higher accuracy of 67.05%, with all errors resulting from incorrect answers.
More specialized questions in identifying conditions and their positions, conditioned on successful abnormality detection, reveal both models’ vulnerabilities to hallucination errors, particularly as they progress through the diagnostic procedure, with Gemini Pro more prone to accepting false conditions and positions. For questions on condition/finding, GPT-4V’s accuracy dropped to 36.9%, with roughly even error distribution between denying ground-truth conditions and accepting hallucinated conditions, while most of the errors of Gemini Pro were from accepting hallucinations. For questions on position, further conditioned on correctly identifying conditions/findings, Gemini Pro had a lower accuracy of 26.4%, with 76.68% of its errors due to accepting hallucinated positions.
@@ 4.2.2 Error Analysis in Procedural Diagnosis
4.2.3 Transferability of Domain Expertise
We conducted a finer-grained analysis to explore whether the model’s expertise in identifying features of a particular organ can be transferred to other imaging modalities. As shown in Table 16, CheXagent, a model trained exclusively on chest X-rays images, performs best in detecting abnormalities and identifying conditions/findings among all seven models when tested on chest X-ray images. We analyzed its performance to explore the transferability of expertise across the rest modalities.
As illustrated in Figure 4, CheXagent achieves significantly higher accuracy in identifying chest-related features compared to other organs, confirming our assumption that the model’s pre-training on chest X-rays enhances its performance on chest images across different modalities. Interestingly, CheXagent also demonstrated higher accuracy in identifying conditions and findings in CT scans and MRIs of the chest, achieving a 3% increase in accuracy for MRIs and a 4% increase for CT scans compared with other organs within the same unseen modality. This indicates that specialized knowledge gained on chest X-rays can be transferred to other imaging modalities of the same organ in a zero-shot manner, highlighting the potential for cross-modality expertise transfer in real-life medical imaging diagnostics.
@@ 5 Conclusion
Evaluating the reliability of LMMs in the medical domain requires robust methods, and ProbMed, our newly introduced dataset, addresses this by incorporating probing evaluation and procedural diagnosis. Our study reveals significant limitations in models like GPT-4V and Gemini Pro, which perform worse than random guessing on specialized diagnostic questions, while CheXagent’s results highlight the importance of domain-specific knowledge. These findings underscore the urgent need for robust evaluation methodologies to ensure the reliability and accuracy of LMMs in medical imaging. Despite the contributions, limitations such as the imbalanced image distribution favoring Chest X-rays (see Table  2) and the lack of open-ended evaluations like report generation remain. The broader impact of our work includes improved diagnostic accuracy and patient care, but also highlights the risks of deploying unreliable models in healthcare. To mitigate these risks, we recommend rigorous testing, continuous performance monitoring, and incorporation of domain-specific expertise. Our work aims to enhance the development of trustworthy AI systems in healthcare, ultimately improving patient outcomes.
@@ A.1 Brain CT Scan
Random Choice acc 25 25 50 35.28 35.01
acc w.o. adv. pair 50 50 / 40.6 40.4
LLaVA-v1-7B acc 25.18 52.59 50 0 0
acc w.o. adv. pair 99.62 87.03 / 92.96 98.94
LLaVA-v1.6-7B acc 10.74 72.22 23.52 0 0.52
acc w.o. adv. pair 99.62 75.18 / 55.18 77.36
LLaVA-Med acc 4.81 10.74 8.82 11.85 3.15
acc w.o. adv. pair 95.55 37.4 / 26.66 92.63
MiniGPT-v2 acc 1.11 92.59 50 17.77 8.42
acc w.o. adv. pair 100 98.88 / 64.81 89.47
CheXagent acc 11.85 0 47.05 12.96 5.26
acc w.o. adv. pair 33.7 0 / 16.66 14.21
GPT-4V acc 94.07 84.07 61.76 37.03 31.05
acc w.o. adv. pair 94.81 84.44 / 55.92 49.47
Gemini Pro acc 84.44 85.18 70.58 34.81 21.05
acc w.o. adv. pair 95.92 91.48 / 45.18 37.36
@@ A.2 Chest CT Scan
Random Choice acc 25 25 50 32.69 33.76
acc w.o. adv. pair 50 50 / 38.89 39.6
LLaVA-v1-7B acc 27.55 46.35 50 0.36 0.23
acc w.o. adv. pair 99.81 99.81 / 96.35 100
LLaVA-v1.6-7B acc 2.73 76.82 50 0.54 0.46
acc w.o. adv. pair 99.08 86.67 / 76.45 87.32
LLaVA-Med acc 5.47 39.78 29.41 14.41 4.37
acc w.o. adv. pair 96.71 90.51 / 35.58 95.62
MiniGPT-v2 acc 0.54 53.28 50 10.21 3.22
acc w.o. adv. pair 99.45 96.53 / 77.55 90.09
CheXagent acc 6.75 72.99 50 18.61 7.83
acc w.o. adv. pair 19.7 99.08 / 35.76 30.41
GPT-4V acc 97.07 72.94 67.64 32.9 20.78
acc w.o. adv. pair 97.62 93.96 / 47.34 39.72
Gemini Pro acc 95.62 58.21 82.35 34.48 14.28
acc w.o. adv. pair 99.27 97.26 / 60.6 61.05

### v5

@@ 4 Experiments and Analysis
Averaged
Accuracy (%)
@@
LLaVA-v1 62.28 25.42 (-36.84) 55.82 19.30 (-36.51)
LLaVA-v1.6 44.06  8.47 (-35.59) 56.02 24.96 (-31.06)
MiniGPT-v2 66.10 46.61 (-19.49) 59.82 27.67 (-32.14)
LLaVA-Med-v1 43.22 3.38 (-39.83) 52.26 17.90 (-34.35)
LLaVA-Med-v1.5 48.30 15.25 (-33.05) 68.41 40.19 (-28.22)
CheXagent 55.50 21.18 (-34.32) 58.70 30.61 (-28.08)
BiomedGPT 56.35 17.79 (-38.55) 60.14 33.34 (-26.79)
Med-Flamingo 61.01 25.42 (-35.59) 64.13 35.66 (-28.47)
RadFM 67.79 38.98 (-28.81) 67.70 41.00 (-26.70)
Gemini Pro 63.13 44.91 (-18.22) 75.08 55.08 (-20.00)
GPT-4V 58.47 33.89 (-24.57) 75.70 55.28 (-20.42)
GPT-4o 69.91 55.08 (-14.83) 76.31 55.60 (-20.71)
We conducted a systematic evaluation and comprehensive analysis using ProbMed on twelve state-of-the-art LMMs to identify their strengths and weaknesses in imaging diagnostics. Apart from proprietary GPT-4o (OpenAI, 2024), GPT-4V (OpenAI, 2023) and Gemini Pro (Team et al., 2024), we selected nine open-source models spanning across general models including LLaVA-v1 (Liu et al., 2024), LLaVA-v1.6 (Liu et al., 2023a), MiniGPT-v2 (Chen et al., 2023) and specialized models including LLaVA-Med-v1, LLaVA-Med-v1.5  (Li et al., 2024), Med-Flamingo (Moor et al., 2023), BiomedGPT (Zhang et al., 2024), RadFM (Wu et al., 2023b) and CheXagent (Chen et al., 2024). These models were chosen based on their computational cost, efficiency, and inference speed, making them practical for integration into medical practice.
4.1 RQ1: Reliability of Current Med-VQA Evaluation
Adversarial Evaluation in VQA-RAD.
To assess whether current Med-VQA evaluations capture model vulnerabilities, we first introduce adversarial pairs on the VQA-RAD test set (Lau et al., 2018). Because VQA-RAD provides only finalized QA pairs without detailed metadata, adversarial pairs were manually constructed by medical experts for 118 “yes” instances (yielding 236 QA pairs total). As shown in Table 2 (left), despite being based on limited information and scale, these adversarial questions lead to drastic accuracy drops. For example, models such as GPT-4o show a reduction from 69.91% to 55.08% (a 14.83% decrease), highlighting the need for robust evaluation protocols.
Adversarial Evaluation in ProbMed.
The ProbMed dataset systematically incorporates adversarial pairs on scale for all 57k QA pairs. Here, each diagnostic entity is paired with a ground-truth question and a corresponding adversarial question. Table 2 (right) demonstrates a similar significant impact: even the best-performing models experience a minimum 20.00% drop in accuracy (with an average decrease of 27.78% across models) when evaluated under this rigorous scheme. This result emphasizes that high accuracy on standard benchmarks can be misleading without essential adversarial evaluation for uncovering model weaknesses.
Table 3: Categorical and overall accuracy (%) of different models aggregated among all image types in ProbMed (averaging over three runs). The overall accuracy is weighted by the number of questions in each type. The best result in each question category is in-bold, and the second best is underlined.
Models General Question Specialized Question Overall
Modality Organ Abnormality Condition/Finding Position
Random Choice 25.00 25.00 50.00 35.67 36.48 32.13
LLaVA-v1 25.30±1.18 41.92±1.21 50.00±2.01 0.35±0.03 0.14±0.06 19.30±0.18
LLaVA-v1.6 6.95±0.24 80.33±0.34 45.89±0.24 3.67±0.10 1.37±0.17 24.96±0.11
MiniGPT-v2 3.25±0.13 76.95±0.59 50.08±0.84 15.23±0.76 7.96±0.79 27.67±0.25
LLaVA-Med-v1 5.72±0.21 34.36±1.21 38.30±2.83 20.79±0.47 5.22±1.10 17.90±0.38
LLaVA-Med-v1.5 56.14±0.90 67.96±0.08 49.12±0.05 21.91±0.06 11.65±0.03 40.19±0.13
CheXagent 37.25±0.50 33.75±0.17 73.31±0.01 28.52±0.08 7.48±0.06 30.61±0.02
BiomedGPT 60.25±0.27 46.81±0.62 50.31±0.24 14.13±0.90 6.11±0.23 33.34±0.17
Med-Flamingo 44.38±0.20 62.02±0.54 50.00±0.01 26.17±0.13 5.72±0.06 35.66±0.14
RadFM 83.72±0.26 41.04±0.33 60.83±0.32 23.05±0.14 9.10±0.29 41.00±0.19
Gemini Pro 96.47±0.88 75.69±1.89 60.29±1.99 27.93±1.82 18.44±0.77 55.08±0.93
GPT-4V 92.51±1.10 71.73±2.45 53.30±1.90 35.19±1.16 22.40±1.89 55.28±0.98
GPT-4o 97.03±0.34 68.13±1.15 61.79±2.28 29.30±2.55 24.06±1.80 55.60±1.05
@@ 4.2 How Reliable Are LMMs in Medical Diagnosis?
After "correcting" inflated model accuracy by introducing adversarial pairs, we continue to address the second research question. We conducted diagnostic probing ranging from general to specialized diagnostic questions using the ProbMed dataset.
Performance across Diagnostic Questions
@@
Table 3 shows the categorical accuracy of different models aggregated among all image types. While GPT-4o, GPT-4V, and Gemini Pro outperform other models and excel in general tasks such as recognizing image modality and organs, their low performance in specialized tasks like determining the existence of abnormalities and answering fine-grained questions about condition/finding and position highlights a significant gap in their ability to aid in real-life diagnosis.
@@
On more specialized diagnostic questions, even top-performing models like GPT-4o, GPT-4V, and Gemini Pro performed close to random guessing. Their accuracy in identifying conditions and positions was alarmingly low, underscoring their limitations in handling fine-grained medical inquiries. RadFM, LLaVA-Med-v1.5, and Med-Flamingo outperform other specialized models in general questions, yet still struggle with specialized questions. LLaVA-Med-v1.5 achieves much higher accuracy among open-sourced models in identifying conditions/finding and their positions, but still performs around 10% worse than the proprietary models.
@@
Among the open-sourced general-purpose models, MiniGPT-v2 performs the best, surpassing domain-specific models such as LLaVA-Med-v1 in determining positions of condition/finding without domain-specific training. A more detailed breakdown of the performance of different models on different image types across each question type is available in Appendix A. Distribution plots of ground-truth answers and model responses within each question category are available in Appendix F.
Error Analysis in Procedural Diagnosis
An error analysis was conducted on two top-performing models (GPT-4V and Gemini Pro) across three specialized diagnostic question types: Abnormality, Condition/Finding, and Position. The procedural diagnosis is conducted by conditioning performance on later and more challenging questions on models’ success on simple prior ones. For example, model accuracy under the Abnormality category is calculated within those test samples where the model correctly identifies both imaging modality and organ. And the accuracy under the Condition/Finding category is further conditioned on those success samples from prior categories, including Abnormality.
As shown in Table 4, both models show vulnerabilities to hallucination errors, particularly in the later stages of the diagnostic procedure. For example, errors under the Abnormality arise either from incorrect responses or a tendency to over-reject challenging questions. In the Condition/Finding and Position categories, errors are largely due to the acceptance of hallucinated entities. Notably, Gemini Pro is more prone to accepting false conditions and positions, which significantly lowers its strict accuracy in these areas.
@@
Figure 4: Accuracy comparison of LLaVA-v1, LLaVA-v1.6, LLaVA-Med-v1, LLaVA-Med-v1.5, Med-Flamingo, and GPT-4o under three different settings: vanilla (baseline performance), chain-of-thought (CoT) reasoning, and CoT with GPT-4o-generated visual descriptions. All models demonstrate significant performance improvement when visual descriptions from GPT-4o are included, indicating that poor visual understanding is a key factor limiting baseline performance. Chain-of-thought reasoning alone also leads to notable gains in accuracy, particularly in general-purpose models.
4.3 Exploring Model Limitations and Potential Improvements
Impact of Chain-of-Thought Prompting and Visual Understanding
To further investigate the underperformance of open-source models, we conducted an extensive ablation study on LLaVA-v1, LLaVA-v1.6, LLaVA-Med-v1, LLaVA-Med-v1.5, Med-Flamingo, and GPT-4o. In this study, we examined two additional experimental settings: (1) applying a Chain-of-Thought (CoT) approach where models first generate visual descriptions from the image on their own, then utilize the description to augment the question, (2) enhancing the models by providing external visual descriptions generated by GPT-4o in addition to the question.
As shown in Figure 4, employing the chain-of-thought approach alone—without external visual descriptions—resulted in an average accuracy increase of 6.51%. In particular, LLaVA-Med-v1.5’s accuracy improved from 40.19% to 54.55%, closing the gap to within 1.05% of the vanilla GPT-4o model. Interestingly, GPT-4o’s performance decreased by 3.55% when the CoT mechanism was applied, potentially indicating that the model already internally employs its own chain-of-thought process.
Notably, all open-source models exhibited improved performance when augmented with visual descriptions generated by GPT-4o, suggesting that their baseline limitations stem primarily from poor visual comprehension. On average, these models showed an accuracy improvement of 9.44% across all question categories. This observation suggests that poor visual understanding is a major limitation of existing models, and augmenting them with external visual reasoning can lead to notable gains. Detailed performance changes of each model, organized by question category, can be found in Appendix D.
Transferability of Domain Expertise
@@ 4.3 Exploring Model Limitations and Potential Improvements
As shown in Table 14, CheXagent, a model trained exclusively on chest X-ray images, performs best in detecting abnormalities and identifying conditions/findings among all models when tested on chest X-ray images. We conducted a finer-grained analysis to explore whether the model’s expertise in identifying features of a particular organ can be transferred to other imaging modalities.
As illustrated in Figure 5, CheXagent achieves significantly higher accuracy in identifying chest-related features compared to other organs, confirming our assumption that the model’s pre-training on chest X-rays enhances its performance on recognizing chest images across different modalities. Interestingly, CheXagent also demonstrated higher accuracy in identifying conditions and findings in CT scans and MRIs of the chest, achieving a 3% increase in accuracy for MRIs and a 4% increase for CT scans compared with other organs within the same unseen modality. This indicates that specialized knowledge gained on chest X-rays can be transferred to other imaging modalities of the same organ in a zero-shot manner, highlighting the potential for cross-modality expertise transfer in real-life medical imaging diagnostics.
@@ 5 Conclusion
Evaluating the reliability of LMMs in the medical domain requires robust methods, and ProbMed, our newly introduced framework, addresses this by incorporating probing evaluation and procedural diagnosis. Our study reveals significant limitations in models like GPT-4o and Gemini Pro, which perform worse than random guessing on specialized diagnostic questions, while CheXagent’s results highlight the critical importance of domain-specific knowledge. Furthermore, our additional experiments, which introduced Chain-of-Thought reasoning with external visual descriptions generated by GPT-4o, suggested that poor visual understanding is a major limitation of existing models and augmenting them with external visual reasoning can lead to notable gains.
6 Limitations
Despite the contributions, limitations such as the imbalanced image distribution favoring Chest X-rays (see Table  1) and the absence of open-ended evaluations, such as report generation, remain. The broader impact of our work includes the potential for improved diagnostic accuracy and better patient care, but it also highlights the risks of deploying unreliable models in healthcare. We recommend rigorous testing, continuous performance monitoring, and the incorporation of domain-specific expertise to mitigate these risks. Ultimately, our work aims to contribute to the development of trustworthy AI systems in healthcare, advancing diagnostic outcomes and patient safety.
Acknowledgments
This research project has benefited from the Microsoft Accelerate Foundation Models Research (AFMR) grant program.
@@ Acknowledgments
This section shows a detailed model performance breakdown by each (imaging modality, organ) pair under five categories.
@@ A.1 Brain CT Scan
Random Choice Acc. with adv. pairs 25 25 50 35.28 35.01
LLaVA-v1 Acc. with adv. pairs 25.18 52.59 50 0 0
Avg. acc. 62.59 72.22 / 46.57 49.60
LLaVA-v1.6 Acc. with adv. pairs 10.74 72.22 23.52 0 0.52
Avg. acc. 55.37 84.44 / 30.79 41.91
MiniGPT-v2 Acc. with adv. pairs 1.11 92.59 50 17.77 8.42
Avg. acc. 50.55 96.29 / 51.20 54.25
LLaVA-Med-v1 Acc. with adv. pairs 4.81 10.74 8.82 11.85 3.15
Avg. acc. 50.18 33.88 / 40.71 49.78
LLaVA-Med-v1.5 Acc. with adv. pairs 50.37 80.37 44.11 11.85 15.26
Avg. acc. 74.81 89.62 / 52.98 54.83
BiomedGPT Acc. with adv. pairs 24.44 5.18 58.82 14.44 2.63
Avg. acc. 62.03 52.59 / 53.88 35.84
Med-Flamingo Acc. with adv. pairs 3.70 9.62 50 18.14 5.26
Avg. acc. 51.85 47.03 / 50.16 47.85
CheXagent Acc. with adv. pairs 11.85 0 47.05 12.96 5.26
Avg. acc. 40.55 23.88 / 53.00 51.46
GPT-4o Acc. with adv. pairs 94.81 93.70 61.76 35.92 26.31
Avg. acc. 97.22 96.66 / 68.76 64.83
GPT-4V Acc. with adv. pairs 94.07 84.07 61.76 37.03 31.05
Avg. acc. 96.85 91.48 / 67.01 65.00
Gemini Pro Acc. with adv. pairs 84.44 85.18 70.58 34.81 21.05
Avg. acc. 92.03 92.40 / 68.01 60.16
@@ A.2 Chest CT Scan
Random Choice Acc. with adv. pairs 25 25 50 32.69 33.76
LLaVA-v1 Acc. with adv. pairs 27.55 46.35 50 0.36 0.23
Avg. acc. 63.77 73.08 / 48.54 50.11
LLaVA-v1.6 Acc. with adv. pairs 2.73 76.82 50 0.54 0.46
Avg. acc. 51.18 86.58 / 41.42 45.75
MiniGPT-v2 Acc. with adv. pairs 0.54 53.28 50 10.21 3.22
Avg. acc. 50.27 75.82 / 51.11 51.49
LLaVA-Med-v1 Acc. with adv. pairs 5.47 39.78 29.41 14.41 4.37
Avg. acc. 51.18 68.06 / 45.50 51.72
LLaVA-Med-v1.5 Acc. with adv. pairs 51.09 61.86 41.17 14.78 9.21
Avg. acc. 75.54 80.10 / 52.60 54.64
BiomedGPT Acc. with adv. pairs 15.51 2.91 52.94 7.11 2.30
Avg. acc. 56.93 50.63 / 50.93 34.65
Med-Flamingo Acc. with adv. pairs 22.26 70.98 50 19.16 7.14
Avg. acc. 60.31 85.49 / 51.11 48.89
CheXagent Acc. with adv. pairs 6.75 72.99 50 18.61 7.83
Avg. acc. 32.93 86.49 / 56.80 51.55
GPT-4o Acc. with adv. pairs 97.62 65.99 67.64 27.60 19.58
Avg. acc. 98.72 81.90 / 63.54 61.67
GPT-4V Acc. with adv. pairs 97.07 72.94 67.64 32.9 20.78
Avg. acc. 98.44 85.74 / 65.01 59.54
Gemini Pro Acc. with adv. pairs 95.62 58.21 82.35 34.48 14.28
Avg. acc. 97.71 78.37 / 65.62 56.84